OpenStudy (anonymous):

the computer repairman is given 8 computers to fix. he knows that among them are 4 bad video cards and 3 failed hard drives. what is the probability thats the first computer he tries has a bad video card but a healthy hard drive? help ?

OpenStudy (dumbcow):

prob of bad video card = 4/8 = 1/2 prob of failed hard drive = 3/8 --> healthy hard drive = 5/8 prob of both = 1/2*5/8 = 5/16